40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 978     Entry time: Mon Sep 22 18:54:54 2008
Author: Jenne 
Type: Update 
Category: PSL 
Subject: PMC transfer functions with various brick-on-top configurations 
Attached below is a graphical summary of different things that I have tried putting on the PMC to reduce the noise in the loop. The motivation behind these measurements is the current inability here at the 40m to increase the UGF of the PMC. This is part of a broader ISS loop/gain/noise problem that we are having, which is causing Rob's locking efforts to have trouble. (The ISS is next on the to-do list, after we find the best configuration for the PMC, if we are still having problems). Right now, it looks like we are being limited by the gain of the PMC (as mentioned by Rana in elog #968).

Anyhow, Rana and I had noticed that piling heavy things on top of the PMC seemed to reduce the noise. What follows are the transfer functions that I took with the different items on top of the PMC, so that we can compare their effects:
  • Nothing on the PMC (like it used to be)
  • New ~14kg lead brick wrapped in copper foil on top of the PMC
  • A stack of a piece of aluminum, a chunk of steel, and then the lead brick on top of the PMC
  • The lead brick + Rob pushing on top of the PMC

Unfortunately, I need to retake the power spectra in these configurations, but from eye-balling it, as one might expect, pushing on the PMC with a hand added more noise than the nominal nothing-on-PMC configuration.

Also unfortunately, none of these configurations seems to have significantly helped our noise reduction situation. We need a new plan. Rana is currently trying out some other configurations, including just aluminum+brick.

Attached is an open loop gain TF from 100Hz - 100kHz. Below that is a zoomed-in version from 5kHz - 30kHz. As you can see more clearly in the zoomed in version, the notch that Rana put onto the board at ~14.5kHz is working, but we need to make the notch deeper, to catch more of that 14.5kHz peak. We're going to try removing the resistor or reducing it's value in the RLC filter on the board (see elog #906). Also, we see that there is a giant peak at 18.3kHz. This is probably much more limiting to our stability at this point than the 14.5kHz peak. We need to add another filter to take care of this, or find another way to reduce this peak. Note that it is present even when there is no brick on the PMC, so it is not an artifact of the new brick.
Attachment 1: PMC_OLG_100Hz_to_100kHz.png  30 kB  | Hide | Hide all
PMC_OLG_100Hz_to_100kHz.png
Attachment 2: PMC_OLG_5kHz_to_30kHz.png  27 kB  | Hide | Hide all
PMC_OLG_5kHz_to_30kHz.png
ELOG V3.1.3-